# Interface-induced turbulence in viscous binary fluid mixtures

### v1

@@ V Conclusions
We have demonstrated how interfaces in a binary-fluid mixture can disrupt low-$Re$ cellular flows by precipitating instabilities that lead to interface-induced turbulence, the binary-fluid analog of elastic turbulence in fluids with polymer additives [1, 2, 3]. We have explored the transitions from cellular flows, to flows with spatiotemporal crystals, and, eventually, to a NESS with interface-induced turbulence. We have characterised these states via the energy time series $e(t)$, its frequency power spectrum $|\tilde{e}(f)|$, the energy spectrum $E(k)$, the energy budget [Eqs. (12) and 13], and the MSD of Lagrangian tracers [Eqs. (9) and (10)]. The low-$\mathrm{Re}$ interface-induced turbulence that we have uncovered exhibits the following distinctive properties: (a) $|\tilde{e}(f)|$ is significant over a broad range of frequencies $f$; (b) a power-law regime with $E(k)\sim k^{-4.5}$, with a power that is different from its counterpart in 2D fluid turbulence with no friction [45, 46]; (c) a scale-by-scale energy transfer [Eq. (12)] with negligible inertial contribution $T(k)$; (d) an MSD of tracers that crosses over from ballistic to diffusive behaviors, indicating strong mixing. Cellular flows have been used in experimental studies of elastic turbulence [50]; we therefore look forward to experimental confirmations of our predictions for low-$\mathrm{Re}$ interface-induced turbulence in such flows.
@@ Acknowledgements.
We thank K.V. Kiran and S. Mukherjee for valuable discussions. NBP, DV, and RP thank the Indo-French Centre for Applied Mathematics (IFCAM) for support and the Isaac Newton Institute for Mathematical Sciences for support and hospitality during the programme ‘Anti-diffusive dynamics: from sub-cellular to astrophysical scales’ when work on this paper was undertaken (EPSRC grant no EP/R014604/1). This research was supported in part by the International Centre for Theoretical Sciences (ICTS) for the online program - Turbulence: Problems at the Interface of Mathematics and Physics (code: ICTS/TPIMP2020/12). NBP and RP thank the Science and Engineering Research Board (SERB) and the National Supercomputing Mission (NSM), India for support, and the Supercomputer Education and Research Centre (IISc) for computational resources. We thank NEC, India for trial use of the SX-Aurora TSUBASA computer, on which we carried out preliminary DNSs for the 2D CHNS system.
$\mathrm{Ca}]{\hskip 455.24408pt}$
0.010.120.170.190.2(a)(f)(k)(p)(b)(g)(l)(q)(c)(h)(m)(r)(d)(i)(n)(s)(e)(j)(o)(t)
Figure 5: Plots for different Capillary numbers Ca increasing from $0.01$ in the bottom row to $0.2$ in the top row: (a)-(e) The scaled total kinetic energy $e(t)/e_{0}$ versus time $t$, with $e_{0}=U^{2}$, the energy of laminar flow without the second phase, and $U=f_{0}/(\nu k_{f}^{2})$ [see Eq (8) in the main text]. (f)-(j) Plots of the corresponding power spectra $|\tilde{e}(f)|$ of $e(t)/e_{0}$. (k)-(o) Pseudocolor plots of the vorticity $\omega$ overlaid with the $\phi=0$ contour (black lines) at a representative time. (p)-(t) Pseudocolor plots of the energy spectra $\mathcal{E}(k_{x},k_{y})$ at a representative time (other parameters are listed in Table 1).
Appendix A Main definitions
We define the energy spectrum and the shell-averaged energy spectrum as
$\mathcal{E}(\mathbf{k},t)=\left[\hat{\bm{u}}(\mathbf{k},t)\cdot\hat{\bm{u}}(-\mathbf{k},t)\right]$ (14)
and
$E(k,t)=\frac{1}{2}\displaystyle\sum_{k\leq|\bm{k}^{\prime}|<k+1}\left[\hat{\bm{u}}(\mathbf{k}^{\prime},t)\cdot\hat{\bm{u}}(-\mathbf{k}^{\prime},t)\right]\,.$ (15)
In the kinetic-energy-budget equation, the transfer term, the contribution of the interfacial stress, and the energy injection term in the kinetic-energy budget equation are
$T(k,t)\equiv-\mathfrak{R}\left[\displaystyle\sum_{k\leq|\bm{k}^{\prime}|<k+1}[\hat{\bm{u}}(-\bm{k}^{\prime},t)\cdot\bm{P}(\bm{k}^{\prime})\cdot\widehat{(\bm{u}\cdot\nabla\bm{u})}(\bm{k}^{\prime},t)]\right],$ (16)
$S(k,t)\equiv\mathfrak{R}\left[\displaystyle\sum_{k\leq|\bm{k}^{\prime}|<k+1}[\hat{\bm{u}}(-\bm{k}^{\prime},t)\cdot\bm{P}(\bm{k}^{\prime})\cdot\widehat{\mathcal{\bm{S}}^{\phi}}(\bm{k}^{\prime},t)]\right]\,,\;\;$ (17)
and
$F(k,t)\equiv\mathfrak{R}\left[\displaystyle\sum_{k\leq|\bm{k}^{\prime}|<k+1}[\hat{\bm{u}}(-\bm{k}^{\prime},t)\cdot\widehat{f^{u}}(\bm{k}^{\prime},t)]\right]\,.$ (18)
$\mathrm{Ca}]{\hskip 156.49014pt}$
0.160.18(a)(b)(c)(d)(e)(f)(g)(h)
Figure 6: Plots for two illustrative Capillary numbers $\mathrm{Ca}=0.16$ (bottom row) and $\mathrm{Ca}=0.18$ (top row): (a),(e) The scaled total kinetic energy $e(t)/e_{0}$ versus time $t$, with $e_{0}=U^{2}$; (b),(f) power spectra $|\tilde{e}(f)|$ of $e(t)/e_{0}$; (c),(g) pseudocolor plots of the vorticity $\omega$ overlaid with the $\phi=0$ contour (black lines) at a representative time; (d),(h) pseudocolor plots of the 2D energy spectra $\mathcal{E}(k_{x},k_{y})$ at a representative time.
Appendix B Simulation parameters and supplementary analysis
The parameters of the simulation are given in Table 1. Plots of the spatiotemporal properties of the flow are shown in Figs. 5 and 6 for different values of the Capillary number Ca.
    R1     R2     R3     R4     R5     R6     R7     R8     R9     R10
      $\mathrm{Ca}$ 0.01 0.1 0.12 0.15 0.16 0.17 0.18 0.19 0.2 0.6
      $\mathrm{Pe}$ 4 36 42 51 54 58 62 66 72 217
   Nature of state STPOG STPOG STPOG STC STPO STPO STC STC STC STPO
Table 1: The non-dimensional Capillary $\mathrm{Ca}$ and Péclet $\mathrm{Pe}$ numbers for our DNS runs R1-R10 ($\mathrm{Re}=1$, $\mathrm{Cn}=0.006$). We fix the following parameters: $f_{0}=0.1024,\,k_{f}=4,\,\nu=0.04,\,\epsilon=0.03,\,M=6\times 10^{-4},\,$ and $\alpha=0$. The natures of the states [nonequilibrium statistically steady states (NESSs) in the chaotic cases] are given in the last row. STPO: spatially and temporally periodic (i.e., a spatiotemporal crystal); STPOG: STPO but with a grain boundary (G) separating two crystalline parts; STC: Spatiotemporal chaos; TPO: temporal periodic oscillations.
Appendix C Video
  • The video showing the spatiotemporal evolution that corresponds to the pseudocolor plots in Fig. 1 is available at: https://youtu.be/yg28KMcSZQw

### v2

@@ V Conclusions
We have demonstrated how interfaces in a binary-fluid mixture can disrupt low-$Re$ cellular flows by precipitating instabilities that lead to interface-induced turbulence, the binary-fluid analog of elastic turbulence in fluids with polymer additives [1, 2, 3]. We have explored the transitions from cellular flows, to flows with spatiotemporal crystals, and, eventually, to a NESS with interface-induced turbulence. We have characterised these states via the energy time series $e(t)$, its frequency power spectrum $|\tilde{e}(f)|$, the energy spectrum $E(k)$, the energy budget [Eqs. (12) and 13], and the MSD of Lagrangian tracers [Eqs. (9) and (10)]. The low-$\mathrm{Re}$ interface-induced turbulence that we have uncovered exhibits the following distinctive properties: (a) $|\tilde{e}(f)|$ is significant over a broad range of frequencies $f$; (b) a power-law regime with $E(k)\sim k^{-4.5}$, with a power that is different from its counterpart in 2D fluid turbulence with no friction [46, 47]; (c) a scale-by-scale energy transfer [Eq. (12)] with negligible inertial contribution $T(k)$; (d) an MSD of tracers that crosses over from ballistic to diffusive behaviors, indicating strong mixing. Cellular flows have been used in experimental studies of elastic turbulence [51]; we therefore look forward to experimental confirmations of our predictions for low-$\mathrm{Re}$ interface-induced turbulence in such flows. In this regard, we note that in our study the two fluids have same kinematic viscosity. Nevertheless, the turbulent state that we have identified is induced solely by the stresses at the interface between the two fluids. We therefore expect it to persist when the kinematic viscosities of the two fluids differ, even though the range of parameters over which interface-induced turbulence is observed is likely to depend on the viscosity ratio.
@@ Acknowledgements.
We thank K.V. Kiran and S. Mukherjee for valuable discussions. NBP, DV, and RP thank the Indo-French Centre for Applied Mathematics (IFCAM) for support and the Isaac Newton Institute for Mathematical Sciences for support and hospitality during the programme ‘Anti-diffusive dynamics: from sub-cellular to astrophysical scales’ when work on this paper was undertaken (EPSRC grant no EP/R014604/1). This research was supported in part by the International Centre for Theoretical Sciences (ICTS) for the online program - Turbulence: Problems at the Interface of Mathematics and Physics (code: ICTS/TPIMP2020/12). NBP and RP thank the Science and Engineering Research Board (SERB) and the National Supercomputing Mission (NSM), India for support, and the Supercomputer Education and Research Centre (IISc) for computational resources. We thank NEC, India for trial use of the SX-Aurora TSUBASA computer, on which we carried out preliminary DNSs for the 2D CHNS system. DV acknowledges the support of the Indo–French Centre for the Promotion of Advanced Scientific Research (IFCPAR/CEFIPRA) (project no. 6704-1).